barry invested $2000, part at 8% annum and the rest at 10% annum. After one year the total interest earned was $190. How much did he invest at each rate?
please help, i have my exam tomorrow and i lost my note on this,

Answer by mananth(16949)

amount invested @ 8 % be x
amount invested @10% will be 2000-x
..
interest equation
0.08x+0.1(2000-x)=190
0.08x + 200-0.1x=190
-0.02x=-10
x=-10/-0.02
x=500
$500 at 8%
balance $ 1500 @ 10%
..
CHECK
1500@10% = 150
500@8%=40
TOtal = 190